### v1

# AlloyASG: Alloy Predicate Code Representation as a Compact Structurally Balanced Graph

Abstract.
In the program analysis and automated bug-fixing fields, it is common to create an abstract interpretation of a program’s source code as an Abstract Syntax Tree (AST), which enables programs written in a high-level language to have various static and dynamic analyses applied. However, ASTs suffer from exponential growth in their data size due to the limitation that ASTs will often have identical nodes separately listed in the tree. To address this issue, we introduce a novel code representation schema, Complex Structurally Balanced Abstract Semantic Graph (CSBASG), which represents code as a complex-weighted directed graph that lists a semantic element as a node in the graph and ensures its structural balance for almost finitely enumerable code segments, such as the modeling language Alloy. Our experiment ensures that CSBASG provides a one-on-one correspondence of Alloy predicates to complex-weighted graphs. We evaluate the effectiveness and efficiency of our CSBASG representation for Alloy models and identify future applications of CSBASG for Alloy code generation and automated repair.
Alloy, Abstract Semantic Graphs, Program Representation, Automatic fixing, Structural balance, Complex-weighted graphs
copyright: none
1. Introduction
Alloy (alloy, 16) is a declarative, relational logic-based modeling language for describing structural and dynamic properties of a system’s design. Since Alloy is used to verify software system designs (Zave15, 40, 5, 34, 8), and to perform various forms of analyses over the corresponding implementation, including deep static checking (JacksonVaziri00Bugs, 17, 10), systematic testing (MarinovKhurshid01TestEra, 22), data structure repair (ZaeemKhurshidECOOP2010, 39), automated debugging (GopinathETALTACAS2011, 12) and to synthesize security attacks (websecurity, 2, 23, 29), the consequences could be catastrophic if an Alloy model is faulty yet successfully compiles. Several iterative search methods based on ASTs have already been implemented for Alloy model automation and repair (amra, 32, 14, 42, 7). However, any iterative search methods for mutants or AST elemental changes that fix the Alloy code segments require expansive storage and have intractable search spaces due to the exponential growth of the code segment length, limiting their computational efficiencies.
In recent years, there have been several bodies of work that focus on automatically repairing programming code using machine learning techniques (mlfix1, 30, 20, 13, 3, 38, 15, 28). Such works provide valuable inspiration to us since we have a dataset of real-world fixing pairs consisting of faulty predicates written by the students and their correct counterparts (alloy4fun, 21), effectively leading to a set of pairs of inputs and ground truths. Nonetheless, almost all the sequence-based algorithms treat the code segments as natural language sequences with a black-box function trained for faulty code as input and corrected code as output. These approaches ignore the strict logic of the programming languages, severely hindering the explainability of the process.
To create a machine learning algorithm that is aware of the logic of the programming language we are generating code in, we need a logical structure-aware, vectorized, machine-and-human-readable, and computationally efficient code representation that improves upon the AST adjacency matrix representation that grows exponentially in size. Some of the representation learning methods (code2seq, 4, 25) train functions that output vectors from the AST, yet the nature of their dependency on the control flow paths in the AST of those general-purpose languages renders their methodologies implausible for a specification language such as Alloy, which is declarative and lacks control flow. Other methods (astnn, 41, 33, 35) try to focus on the raw AST for a universal representation, yet those algorithms still suffer from rapid growth in the size of the AST. Therefore, this paper aims to create a structure that generates a compact representation of the code segment yet still gives the same complete information as the raw AST. Recent work created an Abstract Semantic Graph (ASG) (DuffyASG, 9), which is close to meeting our data representation needs, but the original designation lacks a numerical representation that provides an adjacency or Laplacian matrix that could be used in popular graph machine learning algorithms (gcn1, 24, 11, 18, 31, 27). Therefore, in this paper, we introduce a matrix representation of the ASG with the following objectives:
  1. (1)
    one-on-one correspondence between the source code, the AST, and the graph matrix;
  2. (2)
    better performance in compactness and possible space-saving than the raw AST representation;
  3. (3)
    allowing a direct comparison of the difference between the pair of code segments before and after manual fixing.
Such goals could be impractical for general-purpose, high-level languages such as C or Java, as most languages have a complicated type system and data structures that are not finite or even integer-enumerable. Despite that, for a simple, declarative, almost finitely enumerable language such as Alloy, we could create a Complex-weighted Structurally Balanced Abstract Semantic Graph (CSBASG) that translates the code segments into compact graphs for the subset covering most of the possible combinations.
One of the advantages of the graph is the uniqueness of the semantics. For two semantically different terms in an AST, e.g., a variable or a unary or binary operator, there will only be one corresponding vertex in the CSBASG. This intuitively represents the code elements and their relationships with a reduced space cost. Besides, for a fixing pair of code segments consisting of the faulty code and its correction, CSBASG could align the modifications by the common node signatures shared by the pair and compare their structural differences.
2. Background
2.1. Abstract Syntax Tree and Abstract Semantic Graph
We use the common notation of the abstract syntax tree (AST) for the raw representation of the syntactic structure of the Alloy source code (ast, 6). We follow the formal definitions from (autoinference, 19):
Definition 0 (Context-Free Grammar).
(autoinference, 19) A context-free grammar (CFG) $G\equiv(N,\Sigma,R,s)$ is a tuple where $N$ is the set of nonterminals, $\Sigma$ is the set of terminals, $R$ is the set of production rules $a\to b_{1}b_{2}...b_{N}$ where $a\in N$ is a nonterminal and $b_{i}\in N\cup\Sigma$, and $s\in N$ is the starting nonterminal of the grammar. The language of $G$ is the set of strings from the starting non-terminal.
Definition 0 (AST).
(autoinference, 19) Given a context-free grammar (CFG) $G=(N,\Sigma,R,s)$, an abstract syntax tree (AST) $T\equiv(G,X,r,\xi,\sigma)$ is a tree with a set of nodes $X$, a root node $r\in X$, a mapping $\xi:X\to\mathcal{P}(X)$ from each of the nodes to a subset of nodes as its children and each child set $\forall x\in X:\xi(x)$ is either empty or well-ordered, and $\sigma:X\to(N\cup\Sigma)$ maps each node to its corresponding label as a terminal or nonterminal word in the CFG.
According to (DuffyASG, 9), an abstract semantic graph (ASG) resolves the branches defining the properties of a node in the AST to its original definitions. But here, to create a more compact form of ASG specifically for Alloy, we use a more radical definition that combines every syntactically and semantically equivalent node as below:
Definition 0 (ASG).
Given a CFG $G=(N,\Sigma,R,s)$, an abstract semantic graph (ASG) $\mathcal{G}\equiv(G,\mathcal{V},\mathcal{E},r)$ where $\mathcal{V}\subseteq N\cup\Sigma$ is a set of nonterminal and terminal words as nodes (vertices) of the graph, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}\times\mathbb{C}$ is the set of edges between the words with the edge weight between nodes $v_{i},v_{j}\in\mathcal{V}$ be $w_{ij}\in\mathbb{C}$, and $r$ is the root of the corresponding AST of the ASG, i.e. the first node being visited.
Note that since $\mathcal{E}$ is an arbitrary set of edges in forms of $(v_{i},v_{j},w_{ij})$, an AST or empirical construction of ASG could have more than one set of $\{w_{ij}\}$ to represent. If the weight function $w:\mathcal{V}\times\mathcal{V}\to\mathbb{C}$ ensures that $w_{ij}\neq 0$ if and only if there exists an edge from $v_{i}$ to $v_{j}$, and $w_{ij}\neq w_{ik}$ for any $j\neq k$ be children of $i$, either at the same order of execution (children under the same node in the corresponding AST) or at the different order (if a nonterminal presents more than once in the AST and each time comes with a set of children). There could also be multiple AST links connecting two ASG nodes, and we will discuss how we form a weight value for this case in Section 3.
2.2. Complex-weighted Graph and its Structural Balance
Since Definition 3 gives a complex-weighted graph and our requirement for a vectorization of such ASG, the construction of the graph requires the edge values specifying the relationships between any of the two nodes. So we borrow the structural balance concept from (sbg, 36):
Definition 0 (Complex-Weighted Structural Balanced Graph).
(sbg, 36) For a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, let $a_{ij}=w_{ji}$ be the entries of the adjacency matrix $A$ where $|a_{ij}|>0$ indicates there exists an edge between $v_{j}$ and $v_{i}$, then the graph $\mathcal{G}$ is said to be structurally balanced if all the entries of its adjacency matrix $A=[a_{ij}]^{N\times N}$ satisfies $a_{ij}\equiv|a_{ij}|\angle\theta_{ij}=|a_{ij}|\angle(\theta_{i}-\theta_{j})$, where $\theta_{1},...,\theta_{N}\in(-\pi,\pi)$ is called the signatures of the nodes $v_{1}...v_{N}$, respectively.
In the ASG context, intuitively, the signatures of the nodes $\theta_{1},...,\theta_{n}$ could encode the syntactic and semantic properties of each terminal or nonterminal node and the magnitudes of each edge $|a_{ij}|_{i,j\in\{1...N\}}$ encodes their relative properties, such as if a node is the left (1st) or right (2nd) child of a binary operator. We chose this concept for its mathematical solid properties. Let $d_{i}=\sum_{j\in 1}^{N}|a_{ij}|$ which is the in-degree of node $v_{i}$ and let $D=diag(d_{1},...,d_{N})$ be the diagonal matrix giving the in-degrees of each node on their corresponding rows, and define $L=D-A$ as the Laplacian matrix, then we have the lemma below (sbg, 36):
Lemma 0 ().
(sbg, 36, 37)
The following are equivalent:
(1) Complex weighted graph $\mathcal{G}(A)$ is structurally balanced.
(2) Zero is a eigenvalue of $L$ with eigenvector $\bm{\zeta}=[1\angle\theta_{1},\dots,1\angle\theta_{N}]^{T}$.
(3) $D_{\zeta}:=diag(\bm{\zeta})$ such that $\widehat{A}=D_{\zeta}^{-1}AD_{\zeta}$ is nonnegative and $\widehat{A}=[|a_{ij}|]_{N\times N}$.
(4) $D_{\zeta}:=diag(\bm{\zeta})$ such that $\widehat{L}=D_{\zeta}^{-1}LD_{\zeta}$ has a zero eigenvalue with an eigenvector being $\bm{1}$, where $\widehat{L}=D-\widehat{A}$.
Proof.
See proof of Lemma 1 in (sbgcomplete, 37). ∎
Notably, in the notation of (sbg, 36, 37), each column specifies the outer edges of a node, and the row-based adjacency matrix must be transposed before acquiring the Laplacian.
2.3. Alloy Grammar
We consider the Alloy grammar (alloy, 16) before constructing the graph representation. The basic blocks under the overall root of an Alloy model, ModelUnit, consists of a set of declarations, such as signatures as basic types, functions as processes, and predicates as logical arguments. Our dataset Alloy4Fun (alloy4fun, 21) covers only simple predicates without references to external libraries; therefore, creating representations for those code segments included in the dataset is an ideal starting point.
A predicate defines a scope that contains an expression while itself is a solid, invariant root for any fixing pairs. We can easily enumerate all possible nodes under an expression (or formula): the finitely-enumerable types of expressions or formulae, plus variable declarations that only occur under the quantifiers of first-order logic. In this paper, we assign them fixed signatures; yet an adaptive signature system could be used in case of an extension of the CSBASG into either more types of nodes or extra optimizations.
Another property of Alloy is that almost all expression or formula nodes have a finite and fixed number of children, for instance, 1 for unary operators, 2 for binary operators, and 3 for the if-then-else sentences. In these cases, there could be a trivial, polynomial-based implementation of the $|w_{ij}|$ values for the edges between such a node $v_{i}$ and one of its children $v_{j}$. Even so, there are some exceptions, such as a function call (having an indefinite number of parameters), a list expression or formula (for example, the consecutive logical ANDs and ORs for a conjunctive or disjunctive normal form), or a set of variable declarations for the same quantifier. We will handle them on a case-by-case basis in Section III.
3. CSBASG Construction
Intuitively, to create an ASG following Definition 3, we first need to do a pre-order traversal on the AST to identify and combine the semantically identical nodes. We start by constructing an isomorphism between $\mathcal{V}$ and $N\cup\Sigma$, then rewrite each AST link into an ASG edge. The isomorphism is trivial since AST nodes with the same word are combined into the same node in the ASG. Nonetheless, constructing the ASG edge set by the AST relations $\xi$ requires a cautious step: merging two AST nodes requires clarifying the orders of the visit to avoid confusion. A naive approach that deviates a bit from the ASG definition is to use two numbers for the weight values $w_{ij}$ instead of one, that is, for any of the existing link $x_{2}\in\xi(x_{1})$ in the AST, defining $v_{1}=\sigma(x_{1}),v_{2}=\sigma(x_{2})$. Then, we create an edge between $v_{1},v_{2}$ with a weight-pair $(\omega,t)\in\mathbb{N}\times\mathbb{N}$, specifying that $x_{2}$ is the $\omega$-th child of $x_{1}$ in the order defined for $\xi(x_{1})$, and it is the $t$-th time to visit $v_{1}$ in the program logic.
Nevertheless, we expect that with a matrix form of ASG, we could take advantage of their linearity, especially for the potential application of machine learning operations. In addition, a matrix form of the graphs could be more easily compared with respect to multiple implementations of a predicate or other code segments of interest. With the naive construction above, it is challenging to proceed since the mapping from the pair given above to the edge value that serves as an entry in the adjacency matrix is nontrivial. Here, we could define $\mathcal{M}:\mathbb{N}\times\mathbb{N}\to\mathbb{R}$ as mapping such that $\mathcal{M}(\omega,k)$ gives the entry in the matrix form of an ASG if it is the sole AST link in the ASG. The mapping could be defined arbitrarily, provided that it maintains injectivity to avoid the simple, different-edge-same-value confusion (we will call it Type 1 Confusion later). Such forms the real part (magnitude) of the edges in the CSBASG.
Considering all the above, we formally define the complex-valued structurally balanced ASG as a subset of the general ASG constructions that maintain the structural balance.
Definition 0 (Complex-valued Structurally Balanced ASG (CSBASG)).
A Complex-valued Structurally Balanced ASG is a type of ASG $\mathcal{G}\equiv(G,\mathcal{V},\mathcal{E},r)$ where there exists a vector $\{\theta_{1},\theta_{2},...,\theta_{|\mathcal{V}|}\}$, such that the weights in each of the edge $(v_{i},v_{j},w_{ij})\in\mathcal{E}$ satisfy $w_{ij}=|w_{ij}|\angle(\theta_{i}-\theta_{j})$. For each of the node $v_{i}\in\mathcal{V}$, $\theta_{i}\in(-\pi,\pi)$ is called the signature of $v_{i}$.
Algorithm 1 is a high-level process to convert an AST to a CSBASG. It follows the process of an abstract interpretation, where $\vec{t}$ in the recursive part gives a simplified interpretation state tracking the times of node visitations in the pre-order traversal of the AST. In general, with a static signature assignment $\varphi$, the algorithm terminates in $O(|X|)$ since each of the nodes in the AST was visited exactly once and $|\mathcal{V}|<|X|$ strictly. We say an ASG is generated by an AST if the AST was converted to the same ASG with the same signature assignment and magnitude mapping.
1:Input: an AST $T\equiv(G,X,r,\xi,\sigma)$ parsed from the original code of an Alloy predicate, with root node $r$; an injective function $\varphi:N\cup\Sigma\to(-\pi,\pi)$ to map the words to their unique signature, and another injective function $\mathcal{M}:\mathbb{N}\times\mathbb{N}\to\mathbb{R}$ outputs the magnitude of an AST edge given the positional argument of a link and the time has the sourcing node been visited. For the entries, we use $\beta$ as a recursive encoder that calculates the matrix entries iteratively.
2:Output: a CSBASG $\mathcal{G}(A)$ representing $T$ with its adjacency matrix $A$, and the signatures of the nodes $\{\theta_{1},\theta_{2},...,\theta_{|\mathcal{V}|}\}$, each corresponds to a column of $A$.
3:$\vec{\theta}\leftarrow[0]$
4:$nodemap\leftarrow[(r,0)]$
5:for $i\leftarrow 2,...,|X|$ do
6:     $node\leftarrow X_{i}$
7:     $\theta_{i}\leftarrow\varphi(\sigma(node))$
8:     if $\theta_{i}\notin\vec{\theta}$ then
9:         $\vec{\theta}\leftarrow\vec{\theta}::\theta_{i}$
10:     end if
11:     $nodemap\leftarrow nodemap::(node,\theta_{i})$
12:end for
13:$A_{\vec{0}}\leftarrow[0]_{|\mathcal{V}|\times|\mathcal{V}|}$
14:$A\leftarrow$ encode-recursive($r,\vec{0}_{|\mathcal{V}|},A_{\vec{0}}$)
15:return $A,\vec{\theta}$
16:function encode-recursive($r_{t},\vec{t},A_{\vec{t}}$)
17:     $\theta_{t}\leftarrow nodemap.\textsc{get}(r_{t})$
18:     $i\leftarrow\vec{\theta}.\textsc{get-index}(\theta_{t})$
19:     $t\leftarrow\vec{t}.\textsc{get}(i)$
20:     $\vec{t}^{\prime}\leftarrow\vec{t}$
21:     $\vec{t}^{\prime}[i]\leftarrow t+1$
22:     $A_{\vec{t}^{\prime}}\leftarrow A_{\vec{t}}$
23:     for $\omega\leftarrow 1,...,|\xi(r_{t})|$ do
24:         $c\leftarrow\xi(r_{t}).\textsc{get}(\omega)$
25:         $\theta_{c}\leftarrow nodemap.\textsc{get}(c)$
26:         $j\leftarrow\vec{\theta}.\textsc{get-index}(\theta_{c})$
27:         $A_{\vec{t}^{\prime}}[j,i]\leftarrow\beta(\mathcal{M}(\omega,t),A_{\vec{t}}[j,i])\cdot\angle(\theta_{i}-\theta_{j})$
28:         $A_{\vec{t}^{\prime}}\leftarrow\textsc{encode-recursive}(c,\vec{t}^{\prime},A_{\vec{t}^{\prime}})$
29:     end for
30:     return $A_{\vec{t}^{\prime}}$
31:end function
Algorithm 1 Convert AST to CSBASG
By Definition 5, for a given Alloy predicate as an input, define a fixed unique root $r\equiv v_{1}$, which is the predicate itself, and such root node is unique in each predicate. We can intuitively assign $\theta_{1}=\varphi(\sigma(r))\equiv 0$ as a starting point. Besides, we expect the same input parameters for the pairs of predicates to be directly compared, so the overall root node of each predicate only has one child, the local root expression or formula. Therefore, the only work left is to define the signature assignment function $\varphi$ on the subset that could present under a tree rooted by an expression or formula.
Another notable function, $\mathcal{M}$, could be defined in multiple ways if it maintains a one-on-one correspondence between the syntactic positions and the execution or tree-walking order. However, for an AST with multiple syntactically and semantically identical nodes, multiple edges connect them, but they could have different positions $\omega$ and thus different $\mathcal{M}$ values. So, given an AST, we can define the three types of confusion that could happen in an ASG construction:
Definition 0 (Completeness of CSBASG).
Given a class of AST $T\equiv(G,X,r,\xi,\sigma)$, we say a CSBASG construction is complete if the parser $L$ outputs the AST correctly by the generated CSBASG in Algorithm 1; that is, none of the confusions formed in the construction of $\mathcal{M}$ and $\beta$ mappings:
  • If $\mathcal{M}$ is not injective, i.e., for two edges with positional values $\omega_{1}\neq\omega_{2}$ or tree-walking orders $t_{1}\neq t_{2}$, $\mathcal{M}(\omega_{1},t_{1})=\mathcal{M}(\omega_{2},t_{2})$. This situation is a Type 1 Confusion, or Incomplete Edge Mapping Confusion.
  • If an ASG parser $L$ cannot parse a multi-edge in an ASG that corresponds to two AST links that have the same parent and different but syntactically and semantically equal children with different positions, e.g., $(a+b)-(c+d)$ where the central binary operator -- has two syntactically and semantically equal children of binary operator nodes $+$, then the situation causes a Type 2 Confusion or Twin Children Confusion.
  • If an ASG parser $L$ cannot parse a multi-edge in an ASG that corresponds to two AST links connecting semantically and syntactically equivalent pairs of nodes, but the parent nodes are at different positions in the AST, e.g., $(a\land b)\lor(a\land c)$, the ASG node ``"``"`$\land$"` ` ∧ " has two left children of terminal node $a$, then the situation causes a Type 3 Confusion or Twin Parent Confusion.
Here, the goal becomes creating the mappings $\mathcal{M}$ and $\beta$ that ensure the parser $L$ is free of confusion.
Polynomial-based Static CSBASG Encoding
By Definition 6, intuitively, the signature assignment $\varphi$ is not relevant to the completeness of a CSBASG; that function could be arbitrarily defined or even learned. However, the processes directly related to the fidelity of the representation scheme must be pre-defined to ensure the completeness of the CSBASG. A naive implementation could consider the $\omega$ positional values as sequential numbers $1,2,...$ and add them together for the representation; however, if there is a ternary operator $tor$, then $tor(a,a,c)$ will have its children-position map be defined as $\{a:3,c:3\}$, causes confusion that has two children with a position 3 (and therefore, missing positions 1 and 2), causes a Type 2 confusion.
Here, we introduce the polynomial-based, integer-magnitude static CSBASG encoding for a theoretical language: each AST generated from its CFG has a maximum number of children $p$. Then, let $T$ be the maximum number of syntactic and semantically identical nodes in the code segment. For the AST link drawn from a node to its $(\omega+1)$-th child, let $\mathcal{M}(\omega,t)=2^{(t-1)p+\omega}$ and $\beta$ be a simple summation of the $\mathcal{M}$ values. Our construction of the scheme was defined in the order of $O(2^{pT})$. It could be despairing at a glance, yet consider that in many languages without a list expression, $p\leq 5$ could be achieved, and $T$ only increases with the density and size of the code chunk, which raw AST representation also suffers such a polynomial growth.
Lemma 0 ().
The polynomial-based CSBASG magnitude encoding is complete and optimal; there is no complete encoding with the maximum magnitude of the ASG matrix entry less than $O(2^{pT})$.
Proof.
Completeness follows since $1\leq\omega\leq p$ always holds, for each integer $k\in\mathbb{Z}$ such that $\mathcal{M}(\omega,t)=2^{k}$, there is a unique pair of $\omega$ and $t$, there is no Type 1 Confusion. Consider multiple AST links from the syntactically and semantically same parent node to the syntactically and semantically same child, each with $\omega_{1},\omega_{2},...,\omega_{n}$ as their positions, at $t_{1},t_{2},...,t_{n}$-th visit of the parent node. Then, $\sum_{j=1}^{n}2^{(t_{j}-1)p+\omega_{j}}=\sum_{k=0}^{pT}\mathbb{I}\{\exists{j=1,...,n}|k=(t_{j}-1)p+\omega_{j}\}2^{k}$ is a unique integer representation of the magnitude of an adjacency matrix entry in the ASG, so it is also free of Type 2 or Type 3 Confusion, gives the completeness of the representation. Optimality follows since each positive integer has a binary form $k=\sum_{j=0}^{\infty}c_{j}2^{j}$, $c_{j}\in\{0,1\}$, which each $c_{j}=1$ corresponds to an edge with $j=(t-1)p+\omega$. ∎
By Lemma 2, the summation-of-exponentials gives the most compact and accurate solution for the complete enumeration of the graph, and the creation of such a graph also depends on the enumerability of the grammar. However, a representation within the 32- or 64-bit bound could be achieved for less scaled language, such as Alloy. A heuristic generated by representation learning could also be achieved by learning the logarithmic parameters for a more generally purposed language with more syntactic elements and a larger AST.
The Decoding Algorithm
The decoding process is straightforward in theory. Words of the code are stored in correspondence to rows and columns of the matrix; each row or column corresponds to an individual word, and vice versa. Formally, we designate $L:\mathcal{G}\to\mathcal{T}$ as the parser that translates an ASG back to its uniquely corresponding AST. $L$ is therefore a counterpart of Algorithm 1 and bounded by the constructions of $\mathcal{M}$ and $\beta$ functions. For each of the nonzero entries in the ASG adjacency matrix, we can break it into concrete links by reversing $\mathcal{M}$ and $\beta$. For the binary polynomial-based magnitude encoding given above, we take the integer value of the magnitude $|a_{ij}|$ and break it with regard to its binary form: each digit signing $1$ in the binary form corresponds to a concrete AST link. In this case, an ASG node was broken into a set of AST nodes, each with its corresponding outer edges by the visit order. Since the possible edges are all examined once and the size of the entries determines the binary forms, we have a complexity of $O(|\mathcal{E}|\log(\max(w_{i,j})))$ or $O(|\mathcal{V}|^{2}\log(\max(w_{i,j})))$ for the decoding process.
4. Application of CSBASG on Alloy
Figure 1. An example illustrating the compactness and simplicity of CSBASG. The predicate code (upper left) was parsed into its tree form as the AST (upper right), which has 43 nodes, and it could be simplified to the ASG (lower left) with 15 nodes. The table (lower right) is the correspondence of the nodes in the ASG and its description in the code, and so are the signatures defined and multiplied to the polynomial-based encoding values as shown in the graph. Such constructs a CSBASG.
Alloy, a specification language, has plenty of features that construct a friendly scenario for CSBASG to apply. Firstly, as a first-order logic-based declarative language, we could easily construct a finite set of symbols given a model source code set. Besides, there are few infinitely extensible structures in its grammar, which limits the usage of empirical estimates of the maximum nodes under a given structure; furthermore, even the infinitely extensible structures could be assumed with a relatively low count of children. Its simplicity makes an ideal starting point for CSBASG representation.
To begin with, we utilize a parser that classifies the syntactic properties of the code keywords into a limited number of categories. After that, we need to set a $p$ value corresponding to the maximum number of children for each category, for instance, $1$ for unary operators, $2$ for binary operators, $3$ for if-then-else, and a more significant number for the several uncovered infinite categories. This could cause ambiguity in some edge cases. Still, since Alloy was used in modeling the structures within the scope of a category of sets, there will unlikely be any long enumerations over 10 or 15 items, so we use $p=17$ for the categories without a precise, finite maximum children number. While completeness is not guaranteed, we are aiming at solving the absolute most of the cases; while it is possible to break the enumerations by putting the list of items into a linked structure, in most cases, the tradeoff of completeness in the edge cases could deliver more benefits in intuitiveness and compactness of the representation.
Inside each category, the nodes are differentiated by their semantic contents, which are still finitely enumerable since we have access to the set signatures defined within a model file, and there are almost no numerical constants, so it is safe to assume that constants are limited to booleans and small integers. In our approaches, we divided the AST nodes under a paragraph of an expression or formula into 19 categories, plus the zero-signature node signing the local root. Devolved from the syntactic categories are semantic subcategories, and the number of subcategories within a syntactic category ranges from 1 for the simple, invariant nodes to 32, which is the number of possible binary expressions.
Finally, we are required to assign the angular signatures for each node. In our practice, we used a simple encoding by dividing the interval of a circle $(-\pi,\pi)$ into equal-length intervals with $D=\pi/13$ as the length of each interval. For each interval, we assign subticks that divide it into 128 or 65536 sectors, depending on the syntactic category. By the scheme above, we have obtained each AST node’s syntactic-semantic pair $(\Psi,\psi)$ using the results generated from the parser. Then, we map them on the unit circle as their signatures: each syntactic category goes with the beginning of a $\pi/13$-length interval, and each subtick above the interval starting point corresponds to a semantic subcategory. Note that this is just a demonstration of encoding and is only defined for convenience but not learned for optimality, and the angular correspondence could be any bijective mappings.
Figure 1 gives an example of the encoding from an Alloy predicate AST to its corresponding CSBASG using a polynomial-based static scheme. The syntactic-semantic pairs in the table give the properties of each node in the AST and combine the AST nodes with identical semantic and syntactic properties. The angular signature assigned to each node is $\Psi D+\psi\delta$, where $\delta=D/128$ for expression or formula nodes and $\delta=D/65536$ for the relational declarations that would appear below a quantifying expression or formula, since by Alloy syntax multiple variables could be declared under the same relational declaration node.
The complex values could be used to compare two ASGs with shared parts and will be discussed in Section 5.3.
5. Experiment
Figure 2. An example of a pair of student-written mutant predicate (Inv2) and the ground truth (Inv2C). The ASG edges in form $(x,y)$ indicate the $x$-th visit of the sourcing node, and the positional relationship between the nodes is $y$. Black edges are present in both predicates, red edges are present in the mutant only, and blue edges are present in the ground truth only; a fixing process, in theory, removes all red edges and replaces them with blue ones.
5.1. Dataset
Our dataset encompasses 6,601 models from Alloy4Fun(alloy4fun, 21), sourced from 8 different categories of problem, filtered from a larger dataset to ensure each is compliable and runnable. Each model has two predicates under a set of definitions of relations; a student writes one named InvX (or PropX, X is an identifying number) and is potentially faulty as a mutant predicate, and another called InvXC is the ground truth that gives a correct formula. Then we run two predicates, overconstrained and underconstrained, which determine if there are cases that satisfy InvX but not InvXC, or vice versa. If a model has no cases for both overconstrained and underconstrained predicates, then InvX is formally correct. This gives four categories of the models: correct ($InvX=InvXC$), overconstrained only ($InvX\subsetneq InvXC$), underconstrained ($InvXC\subsetneq InvX$), and both overconstrained and underconstrained ($InvX\not\subset InvXC\land InvXC\not\subset InvX$), covers all possible cases.
5.2. Metric 1: Compactness of representation
We evaluate the compactness of the representation by the percentage of nodes in the ASG representation that are less than the raw AST. Overall, we achieved an impressive 26.6% reduction of nodes over the dataset, considering there is no loss of information, and the performance for each category of predicates is given in Table 1.
Dataset #Models Mut.Type MutantCorrect%
lts 112 CORRECT 28.79 30.17
lts 113 UNDER 24.65 40.52
lts 138 BOTH 24.76 35.9
lts 66 UNDER 27.14 34.31
classroom_fol 536 CORRECT 31.04 14.26
classroom_fol 223 UNDER 30.48 16.86
classroom_fol 1115 BOTH 36.48 26.08
classroom_fol 166 UNDER 37.19 28.54
cv_v1 69 CORRECT 32.28 38.39
cv_v1 118 UNDER 36.05 41.48
cv_v1 68 BOTH 32.49 39.97
cv_v1 55 UNDER 31.45 38.05
train 102 CORRECT 24.77 17.4
train 170 UNDER 23.47 25.23
train 277 BOTH 28.18 28.54
train 178 UNDER 31.96 24.53
trash_rl 340 CORRECT 16.45 6.43
trash_rl 150 UNDER 13.42 7.1
trash_rl 267 BOTH 14.85 6.99
trash_rl 63 UNDER 31.16 11.5
cv_v2 34 CORRECT 30.74 39.42
cv_v2 40 UNDER 38.28 43.79
cv_v2 21 BOTH 38.37 43.08
cv_v2 12 UNDER 29.16 36.21
classroom_rl 536 CORRECT 31.04 14.26
classroom_rl 223 UNDER 30.48 16.86
classroom_rl 1115 BOTH 36.48 26.08
classroom_rl 166 UNDER 37.19 28.54
production 45 CORRECT 17.04 20.57
production 25 UNDER 18.84 20.73
production 22 BOTH 16.02 20.41
production 36 UNDER 22.89 22.8
Table 1. Compactness test results of the dataset. Each of the sub-dataset represents a category of problems within Alloy4Fun(alloy4fun, 21), and the student-written mutant predicates are categorized into four mutant types given their coverage of the set indicated by the ground truth: CORRECT, OVERconstrained, UNDERconstrained, or BOTH over- and underconstrained. The percentages given are the ratio of the reduction of nodes in the ASG representation compared to the raw AST.
Note that since the dataset is mostly single predicates designed as after-class practices for college students, the probability of semantic identical nodes in the AST is relatively low, and for a sub-AST with a large scale, the performance would increase accordingly in the single-metric of node reduction. Moreover, the size of the adjacency matrix scales up with $O(N^{2})$ with the number of nodes, so a denser matrix could be achieved with the ASG scheme.
However, this metric should not be confused with saving spaces. Since in an AST the entries in the adjacency matrix are either 0 (indicating no links) or 1 (indicating a direct link), which only occupy 1-bit per entry in an optimized language or computational method, but Lemma 2 indicates that there is no complete encoding without information loss that takes a number smaller than $O(2^{pT})$ which occupies $pT$-bits at least, so space saving is neither guaranteed or indicated by the inference and experiment above.
5.3. Metric 2: A case study of comparability
Comparing a pair of code snippets to train a model for auto-correction could be tedious, especially for the syntactically correct faulty predicates. Comparability of the CSBASG could point out a direction of inquiry by the graph mutations. We have a simple formula to decode its multiedges, given our construction of the CSBASG with the polynomial method.
Since the dataset was written by students who have little experience in Alloy modeling and some of them write models that have few, if not zero, common edges; however, among them, there could also be some manual repairs that could be shown as a process of applying a set of “Atomic Mutations”. We call an atomic mutation either adding or removing an edge in the ASG; a relationship change between two nodes could also be represented as removing the old edge and adding the new one.
Figure 2 gives an example of the decoded CSBASG construction for comparing two predicates in the same declarative environment, the faulty and correct solution to the same problem. The common parts, as reflected by the subgraph characterized with black edges, show the correct declaration of the first two variables by the student given the right part of the graph and the trivial correctness of the fields associated with the variables as elements of the pre-declared sets were shown in the bottom-left corner. However, most of the graph shows significant deviation of logic, as expected for the student-written mutant being both overconstrained and underconstrained. Straightforwardly, each of the red or blue edges in the graph is an atomic mutation between the pair of predicates, the red edge is a link to be broken, and the blue edge is a link to be established in a hypothetical automatic fixing application.
6. Future Works
6.1. Laplacian of CSBASG and Mutant as a Control Operation
At the beginning of this research, we attempted to use the Laplacian matrix to represent the matrix. However, since some self-edges exist in the scope of the Alloy ASG, and a formally defined construction of a Laplacian matrix did not include this situation, a Laplacian construction for a directed graph containing self-edges is yet to be constructed. Nonetheless, technically, the CSBASG still holds its structural balance, and a direction of inquiry of mutants as control operations in a discrete-time system still holds. There are existing works (acikmese2015spectrum, 1, 26) suggesting constructions that could apply to various kinds of graphs with self-loops, which are potentially fitting constructions to apply for the application that regards a repairing process of a pair of code segments as a control system to develop further possible methods of a modification on a software system containing multiple functional code segments to be considered.
6.2. Learning of Encoding and Alloy Code Generation
In this paper, we only attempted to ensure a zero-information-loss encoding for Alloy code segments but still relied on the map of nodes since we have a static, same-distance pre-defined encoding. Nevertheless, in theory, we can stop this reliance by ensuring no different pair of edges have the same angular signature difference.
Consider a vector $\bm{\zeta}=[1\angle\theta_{1},1\angle\theta_{2},...,1\angle\theta_{n}]^{T}$ where $n$ is the number of total nodes, such that, by Lemma 1, the eigenvector of the CSBASG. We could train this vector as an embedding of the nodes in the CSBASG, giving their angular signatures. In a future scenario, such a vector could be trained with some objective function in applications such as Alloy code generation. A common approach could be an optimization problem like
$\min\sum_{i,j=1...n}\mathbb{I}((v_{i},v_{j})\in\mathcal{E})(\theta_{i}-\theta_{j})$
given $\theta_{i}\neq\theta_{j}$ for any $i\neq j$. For a sub-ASG containing solely expressions or formulae with such a vector $\bm{\zeta}$, we could estimate that it could be a high probability for an incomplete ASG to have edges connecting to the most possible nodes that have a lower angular signature distance with the parent node, helping us to fill out the blanks in an incomplete code segment or generate examples within a predefined declarative environment.
6.3. Automatic Fixing by Machine Learning
Our research was motivated by a requirement to take the programming logic into the machine learning models and create a toolset that could be utilized in automatic fixing. However, due to the limitation of our dataset, few models could be fixed by a small number of mutant operations. Nonetheless, since we already have a mechanism that could output the common edges, edges to be removed, and edges to be established, we could begin with a predictive model for the probability of the existence of each edge, as suggested by some classical graph neural networks, but adapting those methods on the directed multigraph with a polynomial encoding is still posing challenges to the future research.
7. Conclusion
This paper introduced a novel graph representation of the programming language as a complex-valued multigraph with a polynomial-based adjacency matrix encoding. It evaluated its applicability on the declarative specification language Alloy. It further evaluated its improvement of compactness on a dataset consisting of multiple Alloy models and the comparability of the Alloy predicates in an attempt to develop an automatic fixing framework. It points out a direction of inquiry in utilizing the representation in multiple software engineering practices utilizing Alloy and possible integration in a machine learning algorithm.